OpenStudy (anonymous):

Marlene is driving a boat of 180 mph for 3 hours from point a to point b. It took her 2 hours to go back from point b to point a. What is the speed of her going back to point a?

OpenStudy (phi):

I guess they want you to think about rate * time = distance (or speed * time = distance) To use this equation, you need to know 2 of the 3 things. First step, what is the distance between point a and point b ? 180 mph for 3 hours from point a to point b use rate * time = distance

OpenStudy (phi):

you have to notice that 180 mph is a speed and 3 hours is a time

OpenStudy (anonymous):

so 180x3=d?

OpenStudy (phi):

yes. and the answer will be in miles

OpenStudy (anonymous):

240 miles is the answer for it?

OpenStudy (phi):

is your calculator broken ?

OpenStudy (anonymous):

oops I mean 540 I hit the wrong key.

OpenStudy (phi):

ok. the distance between a and b is 540 miles. Common sense says it is 540 miles going from b back to a. next, it says It took her 2 hours to go back from point b to point a. What is the speed of her going back to point a? 2 hours to go back from point b to point a which means 2 hours to go 540 miles (from b back to a) can you find the speed using speed* time = distance

OpenStudy (anonymous):

540x2=d 2580?

OpenStudy (anonymous):

I mean 1080 ^-^

OpenStudy (phi):

does the equation say speed times distance ? the way to do it is write down the equation speed* time = distance replace the 2 things you know: time and distance what do you get ?

OpenStudy (anonymous):

ohh ok now I get it 2/2s=540/2 s=270 is that correct?

OpenStudy (phi):

yes, but I would write it speed* time = distance or s * t = d 2 s = 540 then divide both sides by 2 2 s/2 = 540/2 s = 270 I don't like where you typed 2/2s=540/2 that should be 2s/2 = 540/2

OpenStudy (anonymous):

oh ok sorry. Thank You for helping me!!

OpenStudy (phi):

unless you meant (2/2) s = 540/2 which is ok

OpenStudy (anonymous):

ur medal will be with u shortly ^-^

OpenStudy (phi):

now just say rate times time equals distance 3 times to yourself. and you will be able to do these problems...